# Renormalization Group is the Principle behind the Holographic Entropy Cone

### v1

@@ 3.4.3 Projective plane inequalities
The projective plane inequalities take the form Czech:2023xed; Czech:2024rco:
@@
Our steps mirror Section 3.4.2. Certain values of $f(x)_{j}$ persist for all $x$-bits, which are labeled ‘any’ in Table 7. In each row, taking the union over all implications of saturation indicated in (36) gives:
@@
As in the case of the toric inequalities (Section 3.4.2), these results become more transparent when organized using the graphical formalism developed in Czech:2023xed. They again descend from bit strings $x$, which are represented by minimal loops.
@@ 4 Discussion
For a long time, holographic entropy inequalities have been an answer in search of a question. Previous works have pointed out applications to black hole evaporation Czech:2021rxe and quantum erasure correction Czech:2025jnw and, indeed, the material in Section 3 continues and extends the latter line of thought. But the question posed by the Renormalization Group is general and compelling:
@@
In our view, it does so in a rather clever way. The nonnegative quantities defined by the inequalities are UV-finite and therefore exclusively sensitive to physics away from the asymptotic boundary. Moreover, the notion of bulk depth is adjustable—as indeed it should be in a system described by a conformal field theory. We explained in Section 2.1 that UV-finiteness—manifested by the inequalities with the property of superbalance He:2020xuo—ensures that the relation between holographic RG and the inequalities accommodates every convention of what we designate as the CFT infrared.
The most important finding in this paper is that the saturation of a holographic entropy inequality implies a special structure among entanglement wedges: wedges of larger (infrared) regions, which should ordinarily reach deeper into the bulk than do wedges of smaller (ultraviolet) regions, fail to reach additional depth. By extrapolation, one may speculate that quantum states that violate holographic entropy inequalities can describe bulk geometries, which violate entanglement wedge nesting: ultraviolet wedges reaching deeper than infrared wedges? Engineering such a circumstance requires violating the null energy condition maximin and amounts to constructing a traversable wormhole Gao:2016bin. Can holographic entropy inequalities detect traversable wormholes?
In Section 3, we showed that the saturation of inequalities in the infinite families entails more detailed consequences, which refine and improve result (7). These consequences can be understood both as statements about holographic RG and about access structures of quantum erasure correcting codes Czech:2025jnw. This highlights a relationship between holographic RG and erasure correction, which is obvious but seldom emphasized: running the Renormalization Group improves the protection of logical data. Holographic entropy inequalities endow this slogan with crisp meaning, which is manifested in statements like (45) and Figure 2.
@@
We believe that the RG-related interpretation of holographic entropy inequalities adds evidence in favor of the conjecture of Bao:2025sjn that all such inequalities can be proved by the contraction method. That is because our results rely strongly on contraction proofs. Assuming the connection with RG is not an accident of geometry, it would be strange to find an inequality, which maintains the relationship with the holographic Renormalization Group in spirit but not in its technical content.
@@
An obvious question concerns the issue of time dependence. The techniques we used—first and foremost, contraction maps—assume that all RT surfaces live on a common spatial slice of the bulk, which is manifestly false for HRT surfaces hrt; maximin. Thus, our results—at least in their present form—are strongly anchored to time reversal symmetry. It is also clear that our assertions about spatial slices $E(Y)$ of entanglement wedges do not lift well to full entanglement wedges ${\rm EW}(Y)$, which are the domains of dependence of the $E(Y)$. For example, statement (7) is not true if we substitute $E(Y)\to{\rm EW}(Y)$ for the simple reason that the domain of dependence of the union is not the union of the domains of dependence.
As emphasized in Grimaldi:2025jad, only the issue of time dependence ties the holographic entropy cone to gravitational dynamics. Thinking along these lines has already proven remarkably fruitful for understanding the structure of holographic entropy inequalities. Thus, it is imperative to try to understand whether and how claim (7) can be lifted to time-dependent settings.
Claims (5) and (6) hold under all circumstances, including time dependence and corrections due to bulk entanglement. But it is not clear if one should anticipate claim (7) to have a generalization, which accommodates bulk entanglement. Areas of quantum extremal surfaces Engelhardt:2014gca can famously violate holographic entropy inequalities. (Additional conditions under which they respect them were studied in sergioqes.) The existence of such a generalization is perhaps the most stringent test of the conceptual significance of result (7).
@@ Appendix A Contraction maps
All known holographic inequalities have been proven using a technique called contraction map Bao:2015bfa. For an inequality with the schematic structure (4), one constructs a map $f:\{0,1\}^{L}\to\{0,1\}^{R}$, where $L$ and $R$ count LHS and RHS terms in the inequality. Here is how and why maps $f$ prove entropic inequalities:
@@
In general, many regions $W(x)$ may be empty. When we view the bulk geometry as an erasure correcting code, definition (64) and subregion duality Czech:2012bh; Dong:2016eik assert that the bulk degrees of freedom in $W(x)$ are recoverable from each $X_{i}^{(x_{i})}$ and protected against the erasure of each $\overline{X_{i}^{(x_{i})}}$. (Note that our definitions imply $\overline{X_{i}^{(0)}}=X_{i}$.) Therefore, each bit string $x$ characterizes an access structure, which the code realizes if and only if $W(x)\neq\emptyset$.
@@
When the LHS of the inequality contains pairs $X_{i}\subset X_{j}$ or $X_{i}\subset\overline{X_{j}}$, the map $f$ need not be defined on all of $\{0,1\}^{L}$, but only on bit strings that avoid (67) Avis:2021xnz; lhsconflicts. Note that this restriction automatically sets $x_{i}=x_{j}$ if $X_{i}=X_{j}$, that is when the inequality features a term with a nontrivial coefficient.
@@ Appendix B A contraction-based proof of the Inclusion result
In contrast to the argument in Section 2.1, here we do not assume the conjecture that all holographic entropy inequalities pass the majorization test of Reference Grimaldi:2025jad. We do assume, however, that the inequality can be proved by the contraction method.
@@
and likewise for the norm on $f(x)$. We say that $x^{\prime\prime}$ is colinear with $x,x^{\prime}$ if (79) holds with equality. Colinearity is a useful concept because it imposes many constraints on contraction maps Bao:2015bfa.
A key fact is that UV-finiteness (superbalance) implies that for any two boundary conditions (71) we have He:2020xuo:
@@ Appendix C Prior results: Inequality saturation constrains erasure correcting codes
We review why saturating a valid entropy inequality amounts to a no-go condition on certain access structures in holographic erasure correcting codes. The original argument appeared in Reference Czech:2025jnw, which also contains multiple examples. As elsewhere in the paper, we assume time reversal symmetry in the bulk.
@@
In Reference Czech:2025tds, we studied the occurrences of nesting violations (83) in contraction proofs of all known holographic entropy inequalities. We found that they occur with high density in proofs of all known holographic entropy inequalities except the monogamy of mutual information. Moreover, in cases where this can be quantified, (83) occur with maximal density.
@@
Option (ii) stipulates that the entanglement entropy of $Y_{j}$ is at a phase transition. In this circumstance, any region contained in $E(Y_{j})$ but not in $E^{\rm can}(Y_{j})$ (or vice versa) would be reconstructible from $Y_{j}$ (respectively: protected against erasure of $Y_{j}$) only marginally, rendering the code vulnerable to the tiniest perturbation of the system. Indeed, at holographic phase transitions entanglement entropies are known to be subject to enhanced corrections Dong:2020iod; Akers:2020pmf. Thus, option (ii) describes a scenario that is both fine tuned (phase transition) and unsuitable for constructing erasure correcting codes. For these reasons, in this paper we literally relegate option (ii) to a footnote [^4] In fact, a more careful treatment of option (ii) is possible. To declare that $E^{\rm can}(Y_{j})\neq E(Y_{j})$, one would have to invoke some other hypothetical criterion for deciding on ‘the correct’ entanglement wedge at a phase transition. For example, one might designate the smallest candidate (the overlap of all candidate wedges whose $S^{\rm can}(Y_{j})$ achieve the minimum) to be ‘the correct’ entanglement wedge. In an appendix of Reference Czech:2025jnw, we showed that consistently applying additional criteria does not change the key conclusion of the main argument being presented. and assert simply:

### v2

@@ 3.4.3 Projective plane inequalities
The projective plane inequalities take the form Czech et al. (2024b, a):
@@
Our steps mirror Section 3.4.2. In contraction maps given in Czech et al. (2024b), certain values of $f(x)_{j}$ persist for all $x$-bits, which appear as ‘any’ in Table 7. In each row, taking the union over all implications of saturation indicated in (36) gives:
@@
As in the case of the toric inequalities (Section 3.4.2), these results become more transparent when organized using the graphical formalism developed in Czech et al. (2024b); Czech and Shuai (2025). They again descend from bit strings $x$, which are represented by minimal loops.
@@ 4 Discussion
For a long time, holographic entropy inequalities have been an answer in search of a question. Previous works have pointed out applications to black hole evaporation Czech and Shuai (2022) and quantum erasure correction Czech et al. (2025) and, indeed, the material in Section 3 continues and extends the latter line of thought. But the question posed by the Renormalization Group is general and compelling:
@@
In our view, it does so in a rather clever way. The nonnegative quantities defined by the inequalities are UV-finite and therefore exclusively sensitive to physics away from the asymptotic boundary. Moreover, the notion of bulk depth is adjustable—as indeed it should be in a system described by a conformal field theory. We explained in Section 2.1 that UV-finiteness—manifested by the inequalities with the property of superbalance He et al. (2020)—ensures that the relation between holographic RG and the inequalities accommodates every convention of what we designate as the CFT infrared.
The most important finding in this paper is that the saturation of a holographic entropy inequality implies a special structure among entanglement wedges: wedges of larger (infrared) regions, which should ordinarily reach deeper into the bulk than do wedges of smaller (ultraviolet) regions, fail to reach additional depth. By extrapolation, one may speculate that quantum states that violate holographic entropy inequalities can describe bulk geometries, which violate entanglement wedge nesting: ultraviolet wedges reaching deeper than infrared wedges? Engineering such a circumstance requires violating the null energy condition Wall (2014) and amounts to constructing a traversable wormhole Gao et al. (2017). Can holographic entropy inequalities detect traversable wormholes?
In Section 3, we showed that the saturation of inequalities in the infinite families entails more detailed consequences, which refine and improve result (7). These consequences can be understood both as statements about holographic RG and about access structures of quantum erasure correcting codes Czech et al. (2025). This highlights a relationship between holographic RG and erasure correction, which is obvious but seldom emphasized: running the Renormalization Group improves the protection of logical data. Holographic entropy inequalities endow this slogan with crisp meaning, which is manifested in statements like (45) and Figure 2.
@@
We believe that the RG-related interpretation of holographic entropy inequalities adds evidence in favor of the conjecture of Bao et al. (2025) that all such inequalities can be proved by the contraction method. That is because our results rely strongly on contraction proofs. Assuming the connection with RG is not an accident of geometry, it would be strange to find an inequality, which maintains the relationship with the holographic Renormalization Group in spirit but not in its technical content.
@@
An obvious question concerns the issue of time dependence. The techniques we used—first and foremost, contraction maps—assume that all RT surfaces live on a common spatial slice of the bulk, which is manifestly false for HRT surfaces Hubeny et al. (2007); Wall (2014). Thus, our results—at least in their present form—are strongly anchored to time reversal symmetry. It is also clear that our assertions about spatial slices $E(Y)$ of entanglement wedges do not lift well to full entanglement wedges ${\rm EW}(Y)$, which are the domains of dependence of the $E(Y)$. For example, statement (7) is not true if we substitute $E(Y)\to{\rm EW}(Y)$ for the simple reason that the domain of dependence of the union is not the union of the domains of dependence.
As emphasized in Grimaldi et al. (2025), only the issue of time dependence ties the holographic entropy cone to gravitational dynamics. Thinking along these lines has already proven remarkably fruitful for understanding the structure of holographic entropy inequalities. Thus, it is imperative to try to understand whether and how claim (7) can be lifted to time-dependent settings.
Claims (5) and (6) hold under all circumstances, including time dependence and corrections due to bulk entanglement. But it is not clear if one should anticipate claim (7) to have a generalization, which accommodates bulk entanglement. Areas of quantum extremal surfaces Engelhardt and Wall (2015) can famously violate holographic entropy inequalities. (Additional conditions under which they respect them were studied in Akers et al. (2021).) The existence of such a generalization is perhaps the most stringent test of the conceptual significance of result (7).
@@ Appendix A Contraction maps
All known holographic inequalities have been proven using a technique called contraction map Bao et al. (2015). For an inequality with the schematic structure (4), one constructs a map $f:\{0,1\}^{L}\to\{0,1\}^{R}$, where $L$ and $R$ count LHS and RHS terms in the inequality. Here is how and why maps $f$ prove entropic inequalities:
@@
In general, many regions $W(x)$ may be empty. When we view the bulk geometry as an erasure correcting code, definition (64) and subregion duality Czech et al. (2012); Dong et al. (2016) assert that the bulk degrees of freedom in $W(x)$ are recoverable from each $X_{i}^{(x_{i})}$ and protected against the erasure of each $\overline{X_{i}^{(x_{i})}}$. (Note that our definitions imply $\overline{X_{i}^{(0)}}=X_{i}$.) Therefore, each bit string $x$ characterizes an access structure, which the code realizes if and only if $W(x)\neq\emptyset$.
@@
When the LHS of the inequality contains pairs $X_{i}\subset X_{j}$ or $X_{i}\subset\overline{X_{j}}$, the map $f$ need not be defined on all of $\{0,1\}^{L}$, but only on bit strings that avoid (67) Avis and Hernández-Cuenca (2023); Li et al. (2022). Note that this restriction automatically sets $x_{i}=x_{j}$ if $X_{i}=X_{j}$, that is when the inequality features a term with a nontrivial coefficient.
@@ Appendix B A contraction-based proof of the Inclusion result
In contrast to the argument in Section 2.1, here we do not assume the conjecture that all holographic entropy inequalities pass the majorization test of Reference Grimaldi et al. (2025). We do assume, however, that the inequality can be proved by the contraction method.
@@
and likewise for the norm on $f(x)$. We say that $x^{\prime\prime}$ is colinear with $x,x^{\prime}$ if (79) holds with equality. Colinearity is a useful concept because it imposes many constraints on contraction maps Bao et al. (2015).
A key fact is that UV-finiteness (superbalance) implies that for any two boundary conditions (71) we have He et al. (2020):
@@ Appendix C Prior results: Inequality saturation constrains erasure correcting codes
We review why saturating a valid entropy inequality amounts to a no-go condition on certain access structures in holographic erasure correcting codes. The original argument appeared in Reference Czech et al. (2025), which also contains multiple examples. As elsewhere in the paper, we assume time reversal symmetry in the bulk.
@@
In Reference Czech and Shuai (2025), we studied the occurrences of nesting violations (83) in contraction proofs of all known holographic entropy inequalities. We found that they occur with high density in proofs of all known holographic entropy inequalities except the monogamy of mutual information. Moreover, in cases where this can be quantified, (83) occur with maximal density.
@@
Option (ii) stipulates that the entanglement entropy of $Y_{j}$ is at a phase transition. In this circumstance, any region contained in $E(Y_{j})$ but not in $E^{\rm can}(Y_{j})$ (or vice versa) would be reconstructible from $Y_{j}$ (respectively: protected against erasure of $Y_{j}$) only marginally, rendering the code vulnerable to the tiniest perturbation of the system. Indeed, at holographic phase transitions entanglement entropies are known to be subject to enhanced corrections Dong and Wang (2020); Akers and Penington (2021). Thus, option (ii) describes a scenario that is both fine tuned (phase transition) and unsuitable for constructing erasure correcting codes. For these reasons, in this paper we literally relegate option (ii) to a footnote [^4] In fact, a more careful treatment of option (ii) is possible. To declare that $E^{\rm can}(Y_{j})\neq E(Y_{j})$, one would have to invoke some other hypothetical criterion for deciding on ‘the correct’ entanglement wedge at a phase transition. For example, one might designate the smallest candidate (the overlap of all candidate wedges whose $S^{\rm can}(Y_{j})$ achieve the minimum) to be ‘the correct’ entanglement wedge. In an appendix of Reference Czech et al. (2025), we showed that consistently applying additional criteria does not change the key conclusion of the main argument being presented. and assert simply: